OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Part 2: Provide the solution to the system. 2x – y = 5 x + 2y = 15

jhonyy9 (jhonyy9):

multiply first with 2 and subtract second from first

jhonyy9 (jhonyy9):

4x-2y=10 x +2y =15 ---------- 3x 0 = -5 x= -5/3 -5/3 +2y =15 -5+6y=45 6y=50 y=50/6 y=25/3